Fermi National Laboratory
 
Operation Summary for March 30, 2012 - April 2, 2012
 

NuMI reported receiving 9.05E18 protons on target for the period from 3/26/12 to 4/2/12.

Friday March 30

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest experiments, and the Meson FTBF T-992 experiment taking beam.

During the day shift, Muon Ring personnel began a study at 10:21 AM. At 10:57 AM, the NuMI beam permit tripped off when a power supply’s (E:V118) output drifted low; everything reset. At 1:33 PM, a Booster power supply (B:MP03) dropped to half its normal current; a Booster person investigated and found a bad MADC. Controls personnel changed it out. Operations established beam to MTA at 2:15 PM. SeaQuest made several accesses during this shift and a MI person conducted a g-2 (g minus 2) study during this shift. Operations disabled the MTA beam.

During the evening shift, the Muon Ring study ended at 4:04 PM. At 4:09 PM, the 400 MeV area beam permit tripped off due to a bad beam position; it reset. Operations reported problems with a Linac Klystron RF station at 4:10 PM; a Linac person investigated. At 5:47 PM, a Booster power supply (B:VBC1) tripped off on an overcurrent indication: a Booster person investigated and replaced a firing control logic module. T-992 stopped taking beam for the evening at 9:24 PM. Both SeaQuest and T-992 made accesses during this shift.

Saturday March 31

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 4 AM, the NuMI beam permit tripped off for the first of two times in 94 minutes due to V118; everything reset.

During the day shift, at 8:21 AM, the H- Source high voltage tripped off due to a spark; everything reset. Operations established beam to T-992 around 10:30 AM. The NuMI beam permit tripped off at 11 AM due to an MI incomplete RF gap indication, but no RF station had tripped off; everything reset. At 12:34 PM, the NuMI beam permit tripped off due to V118; everything reset. Operations prepared SeaQuest for access at 1:27 PM.

During the evening shift, at 5:40 PM, Operations reported that they could not turn on a Switchyard (SY) power supply (S:V100) after SeaQuest completed its access; Operators managed to reset the supply locally. Operations disabled beam to T-992 at 8:27 PM.

Sunday April 1

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 1:05 AM, the NuMI beam permit tripped off due to V118; everything reset.

During the day shift, Operations established beam to T-992 around 10:40 AM. At 12:12 PM, the NuMI beam permit tripped off for the first of two times in about an hour due to V118; everything reset.

During the evening shift, at 11:05 PM, the NuMI beam permit tripped off due to V118; everything reset. The T-992 and SeaQuest experiments made accesses during this shift.

Monday April 2

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, at 12:53 AM, the NuMI beam permit tripped off for the first of three times in about three hours due to V118; everything reset. The MI-8 transfer line suffered a vacuum burst at 7:31 AM; it recovered after twenty minutes.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of Muon Ring and MI personnel to conduct studies. SeaQuest continues to commission their beam line. MTA personnel will want beam this week.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/02/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory